Fermi National Laboratory


Accelerator Update for March 28, 2005 - March 30, 2005

Monday March 28
The day shift began with Operations monitoring store 4062, with the Antiproton Source (Pbar) preparing to stash antiprotons into the Recycler, with NuMI off due to a problem with their target, with MCenter off due to the stash, and with MiniBooNE taking beam.

The Recycler finished stashing antiprotons at 8:31 AM.

Pbar resumed taking beam at 8:41 AM.

At 9:09 AM, a Booster RF station (BRF16) tripped off due to a power supply overload. It reset okay.

Operations established beam to the Meson MCenter experiment at 9:19 AM.

At 9:59 AM, the Booster vacuum valves closed and caused a power supply (GMPS) and BRF15 to trip off. This happened a number of times. Operators called in an expert who found a bad vacuum gauge. He completed installing a new gauge at 10:35 AM and Pbar resumed stacking.

At about 10:09 AM, an 8.7 earthquake hit Sumatra. About 15 minutes later, around 10:24 AM, the Tevatron (TeV) motion detectors began recording the event; sound travels about 3000 MPH through rock. The TeV did not quench.

During the Booster downtime, a Linac technician began checking Linac RF stations for LCW leaks. By 10:49 AM, he had found and repaired hose leaks on KRF3 and KRF5.

At 10:58 AM, the Linac tech found and began repairs on a KRF2 hose leak.

Linac and Booster work had all finished by 11:24 AM. Operations resumed sending beam to Pbar, MiniBooNE, and MCenter.

At 11:31 AM, BRF13 tripped off and wouldn't reset. An expert investigated and determined that he would need a tunnel access to make repairs. He managed to reset the station around noon, but it tripped off three more times in forty-five minutes.

At 2:20 PM, the Booster vacuum valves closed again. Operators contacted an expert.

A Preaccelerator expert accessed the H- Source to look at the generator at 2:42 PM. This only took a few minutes.

The Booster expert reported at 3:15 PM that the vacuum gauge PG02 was bad, but this device had two gauges, one upstairs and one down in the tunnel. He switched the interlock to read off of the upstairs gauge. They will repair the tunnel gauge when they have an opportunity to make an access.

Pbar resumed stacking at 3:24 PM. MiniBooNE, MCenter, and MTest also resumed taking beam.

At 4:31 PM, BRF16 tripped off. It reset okay.

At 4:57 PM, BRF13 tripped off. It wouldn't reset, so operators bypassed the station.

At 6:10 PM, Operations tried to turn on NuMI, but a power supply (E:V100) kept tripping off. An operator called in an expert who had the supply repaired two hours later.

At 6:23 PM, BRF16 tripped off. It reset okay.

At 6:26 PM, Operations reported problems with Pbar emittances causing stacking efficiency to drop. Later, Operations reported that the emittance problem was due to excessive stacktail power.

NuMI experts began their target scans at 7:46 PM. They finished at 10:28 PM.

At 7:55 PM, BRF6 tripped off and caused the EAPS to trip off. Operations called in two experts who looked at this and BRF13. They managed to reset BRF6 and BRF13.

At 8:49 PM, BRF6 tripped off. It reset okay.

At 21:24 PM, BRF13 tripped off. It reset okay.

Tuesday March 29
The midnight shift began with Operations monitoring store 4062 and the Recycler stash, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

At 1:20 AM, all the Booster vacuum valves closed. Everything reset okay after a minute.

At 4:06 AM, BRF16 tripped off. It reset okay.

At 4:28 AM, BRF17 tripped off. It reset okay.

Operations terminated store 4062 at 5:03 PM and then put the TeV into shot setup. Any shot setup halts beam to Switchyard (SY).

At 5:22 AM, BRF13 tripped off and would not reset. Operators bypassed the station to let it rest for a bit.

Operations established a combination store 4063 at 7:27 AM with an initial luminosity of 93.31E30. Pbar resumed stacking.

MCenter resumed stacking at 7:58 AM.

At 10:08 AM, BRF15 tripped off twice. It reset okay.

Operations established beam to MTest at 10:47 AM.

Operators turned BRF13 back on at 11:15 AM.

Operations put the Recycler into shot setup to stash antiprotons at 12:31 PM.

At 1:12 PM, turned off Booster for about fifteen minutes to allow work on GMPS3. During that time a Booster expert installed an improved ventilating system in a kicker (MKS02).

The Recycler stash ended at 1:22 PM and Pbar resumed stacking.

At 1:25 PM, BRF13 tripped off and would not reset. The station was bypassed.

A Preaccelerator expert reported at 2:20 PM that the I- Source motor generator had failed.

At 2:21 PM, Operations reported that a Booster power supply (B:DOGL3) tripped off and wouldn't reset. Expert investigated and then gave operators a procedure for resetting the device.

Operations reestablished beam to MCenter and MTest at 2:23 PM.

At 2:58 PM, a Booster kicker (MKS01) tripped off. Experts investigated and found a bad connector. They began replacing the connector. The work took about an hour.

At 4 PM, Operations reported that there was no beam on the Pbar target. Operators and Experts investigated and found a kicker problem (D:IKIK).

At 4:35 PM, BRF6 tripped off. It reset okay.

Pbar resumed stacking at 4:41 PM.

At 6:52 PM, Operations reported that the Pbar stacking efficiency had been terrible since the kicker problem. Operations called in an expert.

MTest reported at 7:13 PM that they were finished with beam for the night.

At 7:16 PM, LRF2 tripped off. Operations reset the station and called in an expert.

An expert had D:IKIK repaired by 8:17 PM. Pbar resumed stacking.

At 8:21 PM, BRF16 tripped off for the first of two times in four minutes. It reset okay.

At 8:20 PM, LRF2 tripped off. It reset okay.

At 8:44 PM, BRF16 tripped off and wouldn't reset. Operators bypassed the station and called in an expert.

At 9:19 PM, Operations reported that Booster was down due to a power supply (WAPS). Operations managed to reset the power supply two minutes later.

At 9:20 PM, LRF2 tripped off for the first of two times in fourteen minutes. It reset okay both times.

A Booster expert brought BRF16 back on line at 9:50 PM.

Pbar resumed stacking at 9:56 PM.

At 10:53 PM, BRF13 tripped off for the first of three times in forty minutes. It reset okay each time.

At 11:32 PM, BRF16 tripped off. It reset okay.

At 11:49 PM, BRF4 tripped off. It reset okay.

Wednesday March 30
The midnight shift began with Operations monitoring store 4063 and the stash, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

At 1:01 AM, BRF11 tripped off. It reset okay.

At 1:13 AM, BRF16 tripped off. It reset okay.

At 7 AM, BRF4 tripped off. It reset okay.

The Plan for Wednesday
The plan for today is to terminate store 4063 around 9 AM and then let the Collider Detector at Fermilab and D-Zero have a two hour access into their collision halls. After the access, Operations will establish a new store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 3/30/2005   email Fermilab